OpenStudy (science0229):

I know that the drift speed can be found by v=J/ne where J is the current density and ne is the carrier charge density. My question is that is ne constant regardless of the cross-sectional area?

OpenStudy (michele_laino):

I think no, since "n" is the density of free electrons with respect to the volume of the conductor

OpenStudy (science0229):

Then, would the drift speed be inversely proportional to the cross-sectional area?

OpenStudy (michele_laino):

no, since J is inversely proportional to the cross sectional area too

OpenStudy (michele_laino):

the drift speed is proportional to the distance traveled by our free electrons

OpenStudy (science0229):

Wait. Since drift speed is proportional to the current density, and the current density is inversely proportional to the cross-sectional area, shouldn't the drift speed be inversely proportional to the cross-sectional area?

OpenStudy (michele_laino):

I think that the reasoning is as below: inside a conductor, we can write: \[\Large {\mathbf{J}} = \sigma {\mathbf{E}}\] so what we have to consider as constant is the electric field inside the conductor

OpenStudy (michele_laino):

\sigma is the conductibility of that conductor

OpenStudy (michele_laino):

now we can write J as follows: \[\Large {\mathbf{J}} = Ne{\mathbf{v}}\]

OpenStudy (science0229):

Then, using this, how can I solve this problem? Figure shows a rectangular solid conductor of edge lengths L, 2L, and 3L. A potential difference V is to be applied uniformly between pairs of opposite faces of the conductor. First, V is applied between the left–right faces, then between the top–bottom faces, and then between the front–back faces. Rank those pairs, greatest first, according to the drift speeds of the electrons.

OpenStudy (michele_laino):

from which we can see that J is proportional to a length, since, dimensionally speaking we have: \[v = \frac{J}{{Ne}} = \frac{{Coulomb/{m^2}}}{{\left( {1/{m^3}} \right)Coulomb}} \sim m\]

OpenStudy (science0229):

okay

OpenStudy (michele_laino):

I try to solve the first case, namely V is applied between left-right faces

OpenStudy (science0229):

You meant to say that v is proportional to the length, right?

OpenStudy (michele_laino):

I mean the length of the space traveled by free electrons

OpenStudy (science0229):

alright

OpenStudy (science0229):

But, isn't the dimensions for the current density A/m^2?

OpenStudy (michele_laino):

sorry you are right! \[v = \frac{J}{{Ne}} = \frac{{Coulomb/\left( {s \times {m^2}} \right)}}{{\left( {1/{m^3}} \right)Coulomb}} \sim \frac{m}{s}\]

OpenStudy (michele_laino):

as we can see the drift speed is proportional to the space traveled by free electrons inside an unitary interval of time

OpenStudy (science0229):

okay

OpenStudy (michele_laino):

ok! for your first case, we can write: \[J = \sigma \frac{V}{{3L}}, \Rightarrow nev = \sigma \frac{V}{{3L}}\]

OpenStudy (michele_laino):

now we can suppose to have N free electrons, inside your conductor, so we can write: \[n = \frac{N}{{volume}} = \frac{N}{{L \times 2L \times 3L}}\]

OpenStudy (michele_laino):

so substituting we have: \[\Large v = \sigma \frac{V}{{3L}}\frac{1}{{ne}} = \sigma \frac{V}{{3L}}\frac{{6{L^3}}}{{Ne}} = \sigma \frac{{2V}}{3}\frac{{{L^2}}}{{Ne}}\]

OpenStudy (science0229):

okay

OpenStudy (michele_laino):

or equivalently: \[\Large v = \sigma \frac{V}{3}\frac{A}{{Ne}},\quad {\text{since}}:A = 2{L^2}\]

OpenStudy (michele_laino):

namely, the speed drift is proportional to the cross sectional area

OpenStudy (michele_laino):